OpenStudy (anonymous):

Rock music accounted for 30.5% of the total sales of recorded music in 1999 , and 20.4% of the total sales of recorded music in 2009 . Find the average rate of change of the percentage of sales of rock music from 1999 to 2009 . Round to the nearest hundredth of a percent. Note that the percentage sign is already given in the answer, so you do not need to enter it. The average rate of change is .

OpenStudy (anonymous):

(30.5 - 20.4)/10 = 10.1/10 = 1.01

OpenStudy (anonymous):

i tried that. it says its wrong :(

OpenStudy (anonymous):

that is because your answer should be negative. sales declined over the ten year period

OpenStudy (anonymous):

\[\frac{30.5-20.4}{1999-2009}=\frac{10.1}{-10}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw